theyadoreshayyy:

One link in a chain was made from a cylinder that has a radius of 2 cm and a height of 20 cm. How much plastic coating would be needed to coat the surface of the chain link? Use 3.14 for π. 251.2cm2 314 cm2 345.4 cm2 471 cm2

theyadoreshayyy:

never mind i figured it out, its A

ItzAgar:

Oh ok lol-

AZ:

@theyadoreshayyy wrote:
never mind i figured it out, its A
That's correct because surface area of cylinder = \( 2\pi r h + 2\pi r^2\) however, this plastic coating only covers the round parts of the cylinder and not the bases, so we're only looking to calculate \( 2\pi r h\) (because the plastic coating is continuous and has to go on to cover the next link in the chain, so it cannot cover the the circular base the radius is 2 the height is 20 2 * pi * r * h 2 * 3.14 * 2 * 20 gets you the answer that you got :)